A small sphere of charge +6 x 10^-6C is suspended by a string of negligible mass. A charge of -9 x 10^-6 C is placed directly to the right of the sphere and 0.22m away from it. The string is deflected 5° from the vertical. Find the tension in the string.

1
Expert's answer
2022-03-14T13:13:09-0400

Answer

Tension is given

"F=\frac{kqq'}{r^2}\\=\frac{9*10^9*6*10^{-6}*9*10^{-6}}{0.22*0.22}\\=10.04N"

"W=mg=1*9.8=9.8N"

So tension

"T=\sqrt{10.04^2+9.8^2}\\=13.90N"